OpenStudy (hannahc234):

What is the degree of the monomial 3x^2y^3 A. 2 B. 3 C. 5 D. 6 What is the simplified form of 8b^3c^2 + 4b^3c^2 A. 12bc B. 12b^3c^2 C. 12b^6c^4 D. 12b^9c^4

OpenStudy (anonymous):

\[ 3x^2y^3\] add the exponents to find the degree

OpenStudy (anonymous):

\[ 8b^3c^2 + 4b^3c^2 \] 8 apples plus 4 apples is 12 apples

OpenStudy (anonymous):

also \[8\diamondsuit + 4\diamondsuit = 12\diamondsuit\]

OpenStudy (anonymous):

you got that?

OpenStudy (hannahc234):

Yes.

OpenStudy (anonymous):

k good

OpenStudy (hannahc234):

So correct me if I'm wrong but.... The first one is C and the second is D?

OpenStudy (anonymous):

first one is C for sure

OpenStudy (anonymous):

the second one is not D don't mess with the exponents leave them as you see them

OpenStudy (hannahc234):

So it's B?

OpenStudy (anonymous):

you only add the exponents when you are multiplying, not adding just like \[100+300=400\] ie. \[10^2+3\times 10^2=4\times 10^2\] yeah B

OpenStudy (hannahc234):

Thank you so much!!

OpenStudy (anonymous):

yw